OpenStudy (anonymous):

check my work please a rigid container of O2 has a pressure of 340 kPa at a temperature of 713 k what is the pressure at 273 k? show work my answer: [P1 ] [ V1 ] / [ T1 ] = [ P2 ] [ V2 ] / [ T2 ] [ P2 ] = [ P1 ] [ V1 / V2 ] [ T2 / T1 ] [ P2 ] = [ 340 ] [ 1.000 ] [ 273.0 / 713.0 ] = 130 kPa.

OpenStudy (anonymous):

I agree with you, but I would exclude V2 and V1from the explanation since they are equal and write the equation as P1/T1=P2/T2 ---> P2=(P1*T2)/T1 Anyway your result is right, I think.

OpenStudy (usman1995):

good

OpenStudy (usman1995):

i haven't read it whole actually lol